• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name jpchen

    Log entry time 18:16:53 on April15,2008

    Entry number 232242

    keyword=cryotarget cool-down and conditions

    We pretty much completed cool-down of cryotarget. Loop 1 (LH2) is at
    20K, 23.8 psi. Fan speed at 36 Hz. HP heater power at ~50 watt. Ready for beam. LD2 is a little high in temperature (26K) and pressure (28psi). Fan speed is at 24 Hz. Since we are not using it for now, I'll leave it as it is until the ESR supply drop to normal (it is dropping from >14K to now 13.8K, and should be ~13.3K). The ESR flow is <12 g/s. We will keep it as it is. Later when ESR becomes stable,
    we will get back to normal operation condition (LD2 = 23.4K, fan=36 Hz) and ESR flow to 10g/s.



    FIGURE 1